OpenStudy (anonymous):

Find the GCF of 14abc and 28a^2b^2c^3

OpenStudy (zehanz):

You can factor everything to see what's going on:\[14abc=2*7*a*b*c\]\[28a^2b^2c^2=2*2*7*a*a*b*b*c*c\]Now can you see what is the GCF?

OpenStudy (zehanz):

Hint: take everything that is the same in both...

OpenStudy (anonymous):

So the answer is 14?

OpenStudy (zehanz):

14abc, because the common factors are : 2, 7, a, b and c.